OpenStudy (anonymous):

The radius of a circle is estimated to be 14 inches, with a maximum error in measure- ment of ±0.08 inches. Use differentials to estimate the maximum error in calculating the area of the circle using this estimate.

OpenStudy (anonymous):

\[A= \pi r^2\]

OpenStudy (anonymous):

\[\frac{dA}{dr} = 2 \pi r \]

OpenStudy (anonymous):

sub in r=0.08

OpenStudy (anonymous):

\[|\epsilon| \le 2(.08)\pi = 1.6\pi \]

OpenStudy (anonymous):

in inches (whatever unit you measured the radius with).